seat exerts a 540-N 60-kg per rce on the rider. (a) radius 16 m, When at the top, the e bottom of the ride? Assume the g-10 ms son rides in a Ferris wheel of spee supporting the person be when at th e person be srmine the speed of the person. (b) What would the force of the seat
Answer #1

At the top,

N = mg - mv2 /r

540 = 60*10 - 60*v2 / 16

540 = 588 - 3.75v2

-60 = - 3.75v2

v = 4 m/s

At the bottom,

N = mg + mv2 /r

N = 60*10 + 60*16 / 16

N = 600 + 60

N = 660 Newtons.